# First Resolution of Microlensed Images of a Binary-Lens Event

### v1

@@ Appendix A Binning the closure-phase data
where $I(\alpha,\delta)$ is the on-sky target intensity distribution centered on light of sight $(\alpha_{0},\delta_{0})$, and $(u,v)=(\bm{n}_{\rm E},\bm{n}_{\rm N})\cdot\bm{B}_{pq}/\lambda$ is the on-sky projection of baseline vector over wavelength. The closure phase is defined for a three telescope array as $\phi_{pqr}(\lambda)=\arg(\mathcal{V}_{pq}\mathcal{V}_{qr}\mathcal{V}_{rp})$. For a list of closure phases $\{\phi_{k}\}$, the mean closure phase $\bar{\phi}$ is calculated by assigning $\phi_{k}$ to unit complex and summing them up for binning as
@@
For given source coordinate $(u_{x},u_{y})$ and binary configuration $(s,q)$, we extract the image contours used for the magnification calculation in VBBinaryLensing package. The contours are then converted to 2D images using the loop-linking method (Dong et al., 2006), which combines the contour integration and inverse ray-shooting methods. The loop-linking method fills the contour with a uniform grid of points organized in vertical links of points. Our code returns the head $(x_{k},y_{k})$, the length $n_{k}$ of every link, and the spacing of the mesh grid $\delta$. For any baseline coordinate $(u,v)$, the complex visibility is calculated by summing up the contribution from all links (assuming $\Psi=0$) as
@@
An alternative method to compute the visibility is applying the divergence theorem on the closed image contour $\partial D$ with
@@
where $\bm{\hat{n}}$ the unit vector normal to the boundary $\partial D$ at $\bm{r}$. For a closed polygon with vertices $\bm{r}_{k}$, the integration for edge $\bm{r}_{k}\to\bm{r}_{k+1}$ is

### v3

@@ Appendix A Binning the closure-phase data
where $I(\alpha,\delta)$ is the on-sky target intensity distribution centered on the light of sight $(\alpha_{0},\delta_{0})$ and $(u,v)=(\bm{n}_{\rm E},\bm{n}_{\rm N})\cdot\bm{B}_{pq}/\lambda$ is the on-sky projection of the baseline vector over wavelength. The closure phase is defined for a three-telescope array as $\phi_{pqr}(\lambda)=\arg(\mathcal{V}_{pq}\mathcal{V}_{qr}\mathcal{V}_{rp})$. For a list of closure phases $\{\phi_{k}\}$, the mean closure phase $\bar{\phi}$ is calculated by assigning $\phi_{k}$ to the unit complex and summing them up for binning as
@@ Appendix B Image generation and visibility calculation
For a given set of source coordinate $(u_{x},u_{y})$ and binary configuration $(s,q)$, we extract the image contours used for the magnification calculation in VBBinaryLensing package. The contours are then converted to two-dimensional images using the loop-linking method (Dong et al., 2006), which combines the contour integration and inverse-ray-shooting methods. The loop-linking method fills the contour with a uniform grid of points organized in links. At a given step size $\delta$, our code returns the head $(x_{k},y_{k})$ and the number of points $n_{k}$ of the $k$th link, which contains a set of points $(x_{k},y_{k}+j\delta)$ with $j\in\{0,1,2\cdots n_{k}-1\}$. For any baseline coordinate $(u,v)$, the complex visibility is calculated by summing up the contribution from all links (assuming $\Psi=0$) as
@@
An alternative method for computing the visibility is applying the divergence theorem on the closed image contour $\partial D$ with
@@
where $\bm{\hat{n}}$ is the unit vector normal to the boundary $\partial D$ at $\bm{r}$. For a closed polygon with vertices $\bm{r}_{k}$, the integration for edge $\bm{r}_{k}\to\bm{r}_{k+1}$ is